# Joint User Association and Resource Allocation for Multi-Cell Networks with Adaptive Semantic Communication

### 05 Jan 2024 v2

@@ IV-C User Association Subproblem
In this subsection, we aim to find the optimal user association decision $\bm{x}^{*}$. We first assume that each WD is associated with all BSs, i.e. $\mathcal{N}_{m}=\mathcal{N}$ for all $m\in\mathcal{M}$. By using algorithms developed in the previous subsections, we can calculate the optimal RB allocation and scheduling decisions. The resulting total utility is an upper bound for the objective value of the original problem (10). After that, we progressively transition to a user association decision that satisfies constraint (1) through $N$ steps. Specifically, in each step, we select a WD that does not satisfy constraint (1) and associate it with exactly one BS. Our objective is to minimize the decrease in utility at each step. The remaining problem is: (i) how to select a WD at each step and (ii) how to choose the target BS for the selected WD.
@@
We first address the second problem. Suppose we have selected a WD $n$ that is associated with all BSs. For an arbitrary BS $m$, let $(z^{m}_{n^{\prime}})_{n^{\prime}\in\mathcal{N}_{m}}$ be the optimal RB allocation. If WD $n$ is no longer associated with BS $m$, the RBs it occupies will be released and allocated to other WDs. The new optimal RB allocatation is denoted by $(\bar{z}^{m}_{n^{\prime}})_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}$. Then the decrease in utility for BS $m$ is
$\Delta_{m}=u^{*}_{n}(z^{m}_{n})-\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}\left(u^{*}_{n^{\prime}}(\bar{z}^{m}_{n^{\prime}})-u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})\right).$
Recall that our objective is to minimize the decrease in total utility. Therefore, we only need to associate WD $n$ with the BS $m$ that has the greatest $\Delta_{m}$. When BS $m$ is associated with multiple WDs, the difference between $z^{m}_{n^{\prime}}$ and $\bar{z}^{m}_{n^{\prime}}$ is relatively small. Therefore, we can use the derivative of $u^{*}_{n^{\prime}}$ at $z^{m}_{n^{\prime}}$ to estimate the difference between $u^{*}_{n^{\prime}}(\bar{z}^{m}_{n^{\prime}})$ and $u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})$, i.e. $u^{*}_{n^{\prime}}(\bar{z}^{m}_{n^{\prime}})-u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})\approx\frac{\mathop{}\!\mathrm{d}u^{*}_{n^{\prime}}}{\mathop{}\!\mathrm{d}z^{m}_{n^{\prime}}}\times(\bar{z}^{m}_{n^{\prime}}-z^{m}_{n^{\prime}})$. If $z^{m}_{n^{\prime}}$ is continuous, then according to the KKT conditions, the derivative $\frac{\mathop{}\!\mathrm{d}u^{*}_{n^{\prime}}}{\mathop{}\!\mathrm{d}z^{m}_{n^{\prime}}}$ is equal for all $n^{\prime}$. In our problem, although $z^{m}_{n^{\prime}}$ takes discrete values, the values of derivatives should be close so we can regard them as approximately equal. Therefore, we only need to calculate the derivative for one $n^{\prime}$. Since it is not very convenient to calculate the derivatives for discrete variables, we use the marginal gain $\delta_{m}=u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}}+1)-u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})$ to replace the role of derivatives. Then we have
$\Delta_{m}$ $\approx u^{*}_{n}(z^{m}_{n})-\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}\delta_{m}(\bar{z}^{m}_{n^{\prime}}-z^{m}_{n^{\prime}})$
$=u^{*}_{n}(z^{m}_{n})-\delta_{m}z^{m}_{n}$ (21)
where the equality holds because $\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}\bar{z}^{m}_{n^{\prime}}=K_{m}$ and $\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}{z}^{m}_{n^{\prime}}=K_{m}-z^{m}_{n}$.
@@
The next problem is how to select the appropriate WD at each step. As described above, for any given WD $n$, we choose its target BS based on $\Delta_{m}$. However, the value of $\Delta_{m}$ depends on the set of WDs associated with BS $m$, which is $\mathcal{N}_{m}$. As our algorithm progresses, $\mathcal{N}_{m}$ also changes. Therefore, the BS previously chosen for WD $n$ may no longer be the optimal one. We refer to this phenomenon as “regret”. Intuitively, if we process the WDs in an appropriate order, the resulting regret will be relatively less.
As mentioned above, regret occurs when the values of $\Delta_{m}$ change so that the previously maximum $\Delta_{m}$ no longer be the maximum. Notice that the values of $\Delta_{m}$ change in a gradual manner as $\mathcal{N}_{m}$ varies (i.e. $\Delta_{m}$ will not change abruptly), hence we have the following two observations. Firstly, for any given WD $n$, if the $\Delta_{m}$ for one BS $m$ is significantly higher than the rest BSs, then it is highly probable that BS $m$ is still the optimal choice for WD $n$ even if $\mathcal{N}_{m}$ changes. Secondly, if WD $n$ is processed relatively late, then the subsequent changes in $\mathcal{N}_{m}$ are limited. This implies that the changes in the values of $\Delta_{m}$ are also limited, hence it is unlikely to result in regret. Based on the above observations, to minimize regret, we should defer the processing of WDs whose values of $\Delta_{m}$ are close. To measure the degree of difference among $\Delta_{m}$, we define the kurtosis of WD $n$ as the ratio of the maximum $\Delta_{m}$ to the sum of all $\Delta_{m}$, i.e. $\kappa_{n}=\frac{\max_{m}\Delta_{m}}{\sum_{m}\Delta_{m}}$. Clearly, if the values of $\Delta_{m}$ are close, then the value of $\kappa_{n}$ is relatively small. To defer the processing of WDs with small $\kappa_{n}$, we only need to pick the WD with the largest $\kappa_{n}$ in each step.
@@
The complete algorithm is summarized in Algorithm 3, where $\mathcal{N}_{r}$ denotes the set of WDs to be processed. Let $K=\max_{m}K_{m}$ be the maximum number of RBs in all BSs. Based on the result in the previous subsection, the time complexity of line $3$ is $O(MNK)$ or $O(MNK^{2})$, depending on whether the utility function is concave. Since the time complexity of the loop in line $4$-$9$ is only $O(NM)$, the time complexity of the whole algorithm is $O(MN^{2}K)$ for concave utility functions and $O(MN^{2}K^{2})$ for general utility functions.
@@ V Numerical Results
TABLE I: Simulation Parameters
@@
Number of RBs in each BS $K_{m}$ $\mathcal{U}\{25,50,75,100\}$
@@
Noise power at BSs $\sigma^{2}$ $10^{-13}$
Average interference at BSs $I_{m}$ $\mathcal{U}[10^{-13},10^{-12}]$
Figure 2: Utility of each BS under the concave utility function.
Figure 3: Utility of each BS under the general utility function.
Figure 2: Utility of each BS under the concave utility function.
Figure 3: Utility of each BS under the general utility function.
@@
Figure 5: Total utility under different numbers of BSs.
@@
Figure 5: Total utility under different numbers of BSs.
Figure 6: Total utility under different numbers of WDs.
@@
In this section, extensive simulations are conducted to evaluate the performance of our algorithm. We consider a wireless network consisting of $M=5$ BSs and $N=30$ WDs that are randomly located in a $\times 500$m square area. For the channel model, we assume the pathloss is $128.1+37.6\log_{10}Dist$ dB and the shadowing factor is $6$ dB, where $Dist$ is the distance (in km) between WDs and BSs.
According to [19], the relationship between result accuracy $A_{d}$ and communication workload $d$ can be modeled by the following function: $A_{d}=\beta^{a}_{1}(1-\frac{d}{D^{a}})^{\beta^{a}_{2}}+\beta^{a}_{3}$, where $\beta^{a}_{1},\beta^{a}_{2},\beta^{a}_{3}$ are parameters and $D^{a}$ is the maximum data size of the DNN outputs. In our simulation, we set $\beta^{a}_{1}\in\mathcal{U}[-0.75,-0.6],\beta^{a}_{2}\in\mathcal{U}[10,20],\beta^{a}_{3}\in\mathcal{U}[0.9,0.95]$ and $D^{a}\in\mathcal{U}[0.15M,0.25M]$ bits. Intuitively, $\beta^{a}_{3}$ is the result accuracy if all DNN outputs are transmitted to the receiver. Hence, the accuracy discount due to limited data transmission is $A_{d}/\beta^{a}_{3}$. Therefore, we assume the final result accuracy at the receiver side is $A=A_{c}\times A_{d}/\beta^{a}_{3}$. To evaluate our algorithm comprehensively, we consider a concave utility function $u^{a}(c,d)=A$ and a general non-concave function $u^{a}(c,d)=\frac{1}{1-A}$. The values of rest parameters are listed in Table I.
  • Traditional Communication (TC): the raw data is transmitted to the BS without processing.
  • Fixed Semantic Communication (FSC): each WD adopts a fixed semantic communication scheme.
  • Nearest User Association (NUA): the WDs are associated with the nearest BSs.
  • FSC+ARB+NUA (FAN): we use the above three algorithms to produce decisions in each stage respectively.
For TC, we assume the size of raw data satisfies $D_{raw}\in\mathcal{U}[0.4M,0.8M]$ bits. If WD $n$ successfully transmits the raw data under the delay and energy constraints, then the BS can use the maximum DNN model to process the raw data, which leads to result accuracy $\eta^{a}_{2}$. For FSC, the utilized fixed semantic communication scheme is $c=D^{a}/2$ and $d=D^{a}/2$. The data presented in this section are the average of ten repeated experiments with different random seeds.
In Fig. 4 and Fig. 4, we show the utility of each BS under the concave and general utility functions, respectively. In both cases, our algorithm achieves the largest utility at all BSs, except for BS $5$, where NUA outperforms our algorithm. This is because there are more WDs around BS $5$ and they are all associated with BS $5$ in NUA, hence leading to a high utility. In our algorithm, however, we will associate part of these WDs with other BSs to achieve better load balance. As a result, the total utility of all BSs (which is shown in Fig. 4) produced by our algorithm is higher than that of NUA. Another algorithm that has comparable performance is ARB. This is not surprising because the marginal gain of result accuracy drops very quickly with the increase of $d$. Consequently, the marginal gain of utility also drops rapidly with the growth of allocated RBs. This guarantees the average allocation of RBs can achieve close-to-optimal results.
Compared to the previous algorithms, the performance of TC and FSC degrades significantly. The problem in TC and FSC is similar: they both adopt fixed transmission schemes so they are unable to efficiently utilize resources when dealing with heterogeneous users and networks. An interesting observation is that the performance of FAN is even significantly worse than TC. This indicates that adopting semantic communication alone does not necessarily enhance system performance if the resource allocation is not optimized. According to Fig. 4, most algorithms exhibit similar performance under the two different utility functions, except for TC. This is because, under the general utility function, the utility is the reciprocal of one minus accuracy, hence the utility increases very quickly when the accuracy approaches one. This is beneficial for TC as it achieves particularly high accuracy when a WD successfully transmits the raw data to the BS.
To understand the impact of network scale on algorithm performance, Fig. 7 and Fig. 7 demonstrate the total utility of all algorithms under different numbers of BSs and WDs. Due to space limitations, we only present the data under the concave utility function, but the result is similar under the general utility function. From Fig. 7, one can observe that the utility of all algorithms increases with the number of BSs. However, when the number of BSs exceeds 6, the utility gains for Prop, ARB, and NUA are very limited. In contrast, the algorithms that adopt fixed transmission schemes maintain a relatively high growth rate. However, in Fig. 7, the growth trend is exactly the opposite. Specifically, the total utility of Prop, ARB, and NUA grows continuously, indicating their ability to serve more WDs with existing communication resources. The only algorithm showing a decreased utility is FAN. This is because FAN adopts the average allocation of RBs, so the number of RBs allocated to each WD becomes less when there are more WDs. Since FAN also employs a fixed transmission scheme, some WDs are unable to complete the transmission under increasingly less RBs, hence resulting in zero utility.
In addition to network scale, we also investigated the impact of delay and energy constraints on the algorithm results, as shown in Fig. 7 and Fig. 8. The algorithms most sensitive to delay are FSC and FAN, because an increase in delay allows WDs to process data in a more energy-efficient manner, thereby improving the probability of successful transmission. The impact of energy budgets on utility is similar, but overall, its effect is less pronounced compared to delay.

### 09 Mar 2026 v3

@@ IV-C User Association Subproblem
In this subsection, we address the optimization of the user association decision $\bm{x}$. We begin by considering a relaxed setting where each WD is temporarily associated with all available SBSs, i.e. $\mathcal{N}_{m}=\mathcal{N}$ for all $m\in\mathcal{M}$. Under this relaxed association, we apply the algorithms developed in previous subsections to compute the optimal RB allocation and scheduling decisions. The resulting total utility serves as an upper bound for the objective value of the original problem (8).
Subsequently, we iteratively refine the user association decision to satisfy the constraint (1) through a sequence of $N$ steps. In each step, we select one WD that violates the association constraint and assign it to exactly one SBS. Our goal in each iteration is to minimize the resulting decrease in total utility. This process naturally raises two questions: (i) how to select the WD to process at each step, and (ii) how to determine its optimal SBS association.
@@
We first address the second question. Suppose we have selected WD $n^{\prime}$ which is currently associated with all SBSs. For an arbitrary SBS $m$, let $(z^{m}_{n})_{n\in\mathcal{N}_{m}}$ be its optimal RB allocation. If WD $n^{\prime}$ is disassociated from SBS $m$, the RBs it previously occupied can be redistributed to other WDs in $\mathcal{N}_{m}\setminus\{n^{\prime}\}$. The new optimal RB allocation is denoted by $(\bar{z}^{m}_{n})_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}$. Then the resulting decrease in utility for SBS $m$ is
$\Delta_{m}$ $=\sum_{n\in\mathcal{N}_{m}}u^{*}_{n}(z^{m}_{n})-\sum_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}u^{*}_{n}(\bar{z}^{m}_{n})$
$=u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})-\sum_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}\left(u^{*}_{n}(\bar{z}^{m}_{n})-u^{*}_{n}(z^{m}_{n})\right).$
Hence, if we eventually associate WD $n^{\prime}$ with SBS $m^{\prime}$ (which means WD $n^{\prime}$ will disassociate with the rest SBSs), then the decrease in total utility will be $\Delta^{t}=\sum_{m\in\mathcal{M}\setminus\{m^{\prime}\}}\Delta_{m}$. Recall that our objective is to minimize $\Delta^{t}$, thus the optimal $m^{\prime}=\operatorname*{arg\,min}_{m^{\prime}}\Delta^{t}=\operatorname*{arg\,max}_{m^{\prime}}\Delta_{m^{\prime}}$.
Since the differences between $z^{m}_{n}$ and $\bar{z}^{m}_{n}$ are typically small, we can employ linear approximation to estimate the change in utility due to a variation in the RB allocation:
$u^{*}_{n}(\bar{z}^{m}_{n})-u^{*}_{n}(z^{m}_{n})\approx\frac{\mathop{}\!\mathrm{d}u^{*}_{n}}{\mathop{}\!\mathrm{d}z^{m}_{n}}\times(\bar{z}^{m}_{n}-z^{m}_{n}).$
Since $z^{m}_{n}$ is the optimal RB allocation, the Karush-Kuhn-Tucker (KKT) conditions imply that the derivative $\frac{\mathop{}\!\mathrm{d}u^{*}_{n}}{\mathop{}\!\mathrm{d}z^{m}_{n}}$ should be equal for all $n\in\mathcal{N}_{m}$.  [^3] Strictly speaking, the KKT conditions apply only to continuous variables. While $z^{m}_{n}$ takes discrete values, the corresponding derivatives are typically approximately equal in practice. Therefore, it suffices to evaluate the derivative for a single WD. Since calculating derivatives for discrete variables is often inconvenient or ill-defined, we approximate the derivative using the marginal gain:
$\delta_{m}=u^{*}_{n}(z^{m}_{n}+1)-u^{*}_{n}(z^{m}_{n}).$
Consequently, the utility reduction $\Delta_{m}$ can be expressed as:
$\Delta_{m}$ $\approx u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})-\sum_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}\delta_{m}(\bar{z}^{m}_{n}-z^{m}_{n})$
$=u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})-\delta_{m}z^{m}_{n^{\prime}}$ (24)
where the equality holds because $\sum_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}\bar{z}^{m}_{n}=K_{m}$ and $\sum_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}{z}^{m}_{n}=K_{m}-z^{m}_{n^{\prime}}$.
@@
The next problem is how to select the appropriate WD to process at each step. As described above, for any given WD $n$, we choose its target SBS based on $\Delta_{m}$. However, the value of $\Delta_{m}$ depends on the set of WDs associated with SBS $m$, which is $\mathcal{N}_{m}$. As WDs progressively assigned to specific SBSs, the set $\mathcal{N}_{m}$ evolves, potentially altering the values of $\Delta_{m}$. Consequently, a previously optimal SBS for WD $n$ may no longer be optimal later—a phenomenon we term regret.
To mitigate regret, we make the following two observations. First, if the value of $\Delta_{m}$ for one SBS significantly exceeds that of others, it is likely that the SBS remains optimal even as $\mathcal{N}_{m}$ changes. Second, if WD $n$ is processed late in the sequence, then the subsequent changes in $\mathcal{N}_{m}$ are limited. This implies that the variation of $\Delta_{m}$ is also limited, thereby reducing the risk of regret.
Based on these observations, we prioritize WDs for which the values of $\Delta_{m}$ differ most distinctly. To quantify this, we define the kurtosis of WD $n$ as:
$\kappa_{n}=\frac{\max_{m}\Delta_{m}}{\sum_{m}\Delta_{m}}.$
A lower $\kappa_{n}$ indicates that $\Delta_{m}$ values are more uniform across SBSs, implying higher uncertainty in SBS selection. Hence, in each step, we select the WD with the largest $\kappa_{n}$, thereby deferring those with higher risk of regret.
@@
The overall procedure is outlined in Algorithm 3, where $\mathcal{N}_{r}$ denotes the set of WDs to be processed. Let $K=\max_{m}K_{m}$ be the maximum number of RBs in all SBSs. Based on the result in the previous subsection, calculating the optimal RB allocation in line $3$ has a time complexity of $O(MNK)$ for concave utility functions and $O(MNK^{2})$ for general ones. The time complexity of user selection and update process in line 44-99 is $O(NM)$. Thus, the total time complexity of Algorithm 3 is $O(MN^{2}K)$ for concave utility functions and $O(MN^{2}K^{2})$ for general utility functions.
@@ V Numerical Results
TABLE II: Simulation Parameters
@@
Number of RBs in each SBS $K_{m}$ $\mathcal{U}\{10,20,30,40\}$
@@
Noise power at SBSs $\sigma^{2}$ $10^{-13}$
Average interference at SBSs $I_{m}$ $\mathcal{U}[10^{-13},10^{-12}]$
In this section, extensive simulations are conducted to evaluate the performance of our algorithm and numerical results show that our method significantly outperforms the conventional communication paradigm and existing semantic compression approach.
V-A Simulation Setup
We consider a 5G network consisting of $M=10$ SBSs and $N=100$ WDs that are randomly located in a $1000\times 1000m$ square area. For the channel model, we assume the pathloss is $128.1+37.6\log_{10}Dist$ dB and the shadowing factor is $6$ dB, where $Dist$ is the distance (in km) between WDs and SBSs.
According to [46], the relationship between result accuracy $A_{d}$ and communication workload $d$ can be modeled by the following function: $A_{d}=\beta^{a}_{1}(1-\frac{d}{D^{a}})^{\beta^{a}_{2}}+\beta^{a}_{3}$, where $\beta^{a}_{1},\beta^{a}_{2},\beta^{a}_{3}$ are parameters and $D^{a}$ is the maximum data size of the DNN outputs. In our simulation, we set $\beta^{a}_{1}\in\mathcal{U}[-0.75,-0.6],\beta^{a}_{2}\in\mathcal{U}[10,20],\beta^{a}_{3}\in\mathcal{U}[0.9,0.95]$ and $D^{a}\in\mathcal{U}[0.15M,0.25M]$ bits. Intuitively, $\beta^{a}_{3}$ is the result accuracy if all DNN outputs are transmitted to the receiver. Hence, the accuracy discount due to limited data transmission is $A_{d}/\beta^{a}_{3}$. Therefore, we assume the final result accuracy at the receiver side is $A=A_{c}\times A_{d}/\beta^{a}_{3}$. To evaluate our algorithm comprehensively, we consider a concave utility function $u^{a}(c,d)=A$ and a general non-concave function $u^{a}(c,d)=\frac{1}{1-A}$. The values of rest parameters are listed in Table II.
  • Traditional Communication (TC): the raw data is transmitted to the SBS without processing.
  • Fixed Semantic Communication (FSC): each WD adopts a fixed SemCom scheme.
  • Semantic Compression (SC): fixed computation with selective transmission of semantic information.
  • Nearest User Association (NUA): the WDs are associated with the nearest SBSs.
For TC, we assume the size of raw data satisfies $D_{raw}\in\mathcal{U}[0.4M,0.8M]$ bits. If WD $n$ successfully transmits the raw data under the delay and energy constraints, then the SBS can use the maximum DNN model to process the raw data, which leads to result accuracy $\eta^{a}_{2}$. For FSC, the utilized fixed SemCom scheme is $c=C^{a}/2$ and $d=D^{a}/2$. In SC schemes [31, 46, 20], the complete semantic representation is first extracted, resulting in $c=C^{a}$. Subsequently, only the most critical semantic information is transmitted, subject to constraints on communication resources, latency, and energy consumption. To further validate the effectiveness of our design, we introduce two ablation baselines, namely ARB and NUA, which are employed to isolate and highlight the contributions of the proposed RB allocation and user association strategies, respectively. The data presented in this section are the average of $30$ repeated experiments with different random seeds.
Figure 2: Locations of WDs and SBSs.
Figure 3: Utility of each SBS under the concave utility function.
V-B Performance Comparison under a Specific Scenario
To elucidate the strategic differences among the considered algorithms, we first analyze system performance under a representative scenario. Fig. 2 illustrates the randomly generated positions of WDs and SBSs, while Fig. 3 presents the utility achieved by each SBS under the concave utility function. In our simulations, the available wireless resources are highly constrained relative to the volume of raw data. As a result, either none or only a limited fraction of WDs can successfully transmit their raw data to SBSs, leading to degraded system utility under TC. The adoption of SemCom alleviates this issue by substantially reducing the amount of data to be transmitted. Within this framework, SC outperforms FSC owing to its more adaptable transmission strategy. Nevertheless, SC still suffers from fixed computational requirements, resulting in excessive latency and energy consumption.
In contrast, ASC achieves a notable performance gain. Due to the heterogeneity of WDs, equal allocation of RBs across users is inefficient. Consequently, although the ablation baseline ARB achieves higher utility for certain SBSs, the overall network utility remains inferior to that of the proposed approach. A similar observation applies to the NUA strategy, which tends to yield near-optimal performance in regions with sparsely distributed SBSs. However, in areas where SBSs are densely located, systematic user association becomes crucial. For instance, by associating some WDs in proximity to SBSs $1$ and $2$ with SBS $6$, the proposed method further enhances overall network utility.
@@ V-B Performance Comparison under a Specific Scenario
V-C Performance under Different Utility Functions
To provide a comprehensive comparison of the overall performance of the considered algorithms, Fig. 4 depicts their performance under both concave and general utility functions. For the concave utility function, the performance trends of the algorithms are consistent with the case study presented in the previous subsection. Moreover, most algorithms exhibit comparable performance across the two utility functions, with the exception of TC. This divergence arises because, under the general utility function, utility is defined as the reciprocal of one minus accuracy, which increases sharply as accuracy approaches unity. Consequently, TC benefits from this property, as it achieves particularly high utility when a WD successfully transmits the raw data to the SBS. In both utility settings, the proposed algorithm consistently achieves the highest utility, thereby demonstrating its robustness and near-optimality across different classes of utility functions.
Figure 5: Total utility under different numbers of SBSs.
@@ V-C Performance under Different Utility Functions
V-D Impact of Network Scale
To evaluate the influence of network scale on algorithmic performance, Fig. 5 and Fig. 6 present the total utility of all algorithms under varying numbers of SBSs and WDs. For clarity, only results under the concave utility function are shown, as similar trends are observed for the general utility function. As illustrated in Fig. 5, increasing the number of SBSs reduces the average distance between WDs and SBSs, thereby improving channel gains. Moreover, a larger number of SBSs also provides additional RBs. Consequently, the utility achieved by all algorithms increases with the number of SBSs. However, the performance improvement of TC remains marginal, highlighting its inefficiency. Among the ASC-based schemes, ARB exhibits a smaller performance gain compared with the proposed algorithm and NUA, primarily because its average RB allocation strategy fails to fully exploit the additional RBs, thereby diminishing the benefits of network densification.
A similar trend is observed when varying the number of WDs. Given the limited wireless resources, systematic resource allocation becomes critical to achieving higher utility. Nonetheless, since both TC and FSC employ fixed transmission strategies, they lack enough flexibility in resource allocation, resulting in minimal performance gains. For the same reason, although SC adopts an adaptive transmission strategy, its fixed computation mechanism restricts flexibility compared to ASC paradigms. Consequently, its performance improvement is also limited relative to the proposed algorithm, NUA, and ARB.
@@ V-D Impact of Network Scale
V-E Impact of Delay and Energy Constraints
In addition to network scale, we further examine the effects of delay and energy constraints on algorithmic performance, as depicted in Fig. 7 and Fig. 8. Among all schemes, SC exhibits the highest sensitivity to both constraints. This behavior arises because SC employs a fixed computation process that consumes a constant amount of time and energy; thus, any additional delay or energy budget can be entirely utilized for transmitting more semantic information, resulting in substantial marginal gains. Furthermore, the total utility of all algorithms is observed to be more sensitive to delay requirements than to energy budgets. This outcome is not surprising, as the transmitted data volume scales linearly with transmission time but only logarithmically with transmission power. Notice that our algorithm constantly outperforms all benchmarks in all cases, demonstrating its effectiveness under diverse scenarios.